A garden is currently 4 m wide and 7 m long. If the area of the garden is doubled by increasing the width and length by the same number of meters, find the new dimensions of the garden.
Found 3 solutions by scott8148, nyc_function, acoldwell:
Answer by scott8148(6628)

(4 + x)(7 + x) = 2(4 * 7)

28 + 11x + x^2 = 56

x^2 + 11x - 28 = 0

use quadratic formula to find x

Answer by nyc_function(2741)

The area of the original garden is A = 7 * 4 or 28m^2.
The length and width is increased by the same amount of meters.
Let x = meter amount we want
We are also told the area is doubled when we add this unknown number to the length and width.
So, 28 * 2 = 56.
56 = (x + 7)(x + 4)
56 = x^2 + 11x + 28
0 = x^2 + 11x + 28 - 56
0 = x^2 + 11x - 28
All you have to do is factor this quadratic equation using the quadratic formula. When you find x, add your value of x to the length (x + 7) and the width (x + 4).
Can you finish now?

Answer by acoldwell(1)

okay, this is an easy one,
Step 1) double both the length and width. now the length should equal 14m and the width should equal 8m
Step 2) now you can solve for the area, by using Length x Width and you should come up with 112m

if you have any questions , feel free to just ask!
-amber